The Paulson Company's year-end balance sheet is shown below. Its cost of common equity is 18%, its before-tax cost of debt is 9%, and its marginal tax rate is 25%. Assume that the firm's long-term debt sells at par value. The firm’s total debt, which is the sum of the company’s short-term debt and long-term debt, equals $1,150. The firm has 576 shares of common stock outstanding that sell for $4.00 per share.

Assets Liabilities And Equity
Cash $ 120 Accounts payable and accruals $ 10
Accounts receivable 240 Short-term debt 60
Inventories 360 Long-term debt 1,090
Plant and equipment, net 2,160 Common equity 1,720
Total assets $2,880 Total liabilities and equity $2,880

Calculate Paulson's WACC using market-value weights. Do not round intermediate calculations. Round your answer to two decimal places.

%
Answer #1

Value of equity =No of shares outstanding*share price =576*4 =2304

Value of debt =1150

Total value =2304+1150 =3454

After-tax cost of debt =9%*(1-25%) =6.75%

WACC=(2304/3454*18%)+(1150/3454*6.75%) =14.25%